# Nested Graph Pseudo-Label Refinement for Noisy Label Domain Adaptation Learning

### v2

@@ Introduction
Graph Domain Adaptation (GDA) (You et al. 2022; Cai et al. 2024) has emerged as a prominent technique for leveraging labeled graph data from a source domain to enhance learning on an unlabeled target graph domain. Its efficacy has been demonstrated across diverse applications, including temporally-evolved social network analysis (Wang et al. 2021, 2024c; Yao et al. 2023), molecular property prediction (Zhu et al. 2023; Yin et al. 2024c), and protein-protein interaction modeling (Cho, Berger, and Peng 2016; Wang et al. 2024a). The core paradigm typically involves learning domain-invariant node/graph representations that bridge the distributional shift between source and target domains, thus enabling effective inference on the target data.
@@
However, the success of standard GDA methods crucially relies on the accurately labeled source data. In practice, source domain labels are often corrupted by noise arising from annotation errors (Dai, Aggarwal, and Wang 2021; Yuan et al. 2023b), subjective judgments (Platanios, Dubey, and Mitchell 2016; Yin et al. 2024d), or inherent ambiguities in data collection (Chen, Shah, and Kyrillidis 2020; Wang et al. 2024b). This prevalent issue of label noise can severely misguide the learning of domain-invariant representations (Li, Socher, and Hoi 2020; Yin et al. 2024b), leading to suboptimal or even detrimental adaptation performance on the target domain (Yin et al. 2025; Wang et al. 2025). Existing noise label learning methods typically rely on loss function design to mitigate the impact of noisy labels (Han et al. 2018; Natarajan et al. 2013), which selects clean instances for joint training, and robust loss functions (Wei et al. 2020; Li, Socher, and Hoi 2020), which leverage small-loss selection or instance mixture models. While effective in controlled settings, these approaches fall short in the presence of domain shifts. The coexistence of distribution shift and label noise leads to misaligned feature spaces, causing noise-robust losses to erroneously align clean features with noisy targets, thereby amplifying negative transfer (Yu et al. 2020; Yin et al. 2022a). While recent efforts have been made to address GDA under noisy labels (Yuan et al. 2023a; Wang and Yang 2022), they primarily target node classification tasks, leaving a critical gap in addressing graph-level scenarios. Many real-world applications, such as molecular property prediction (Stokes et al. 2020) and social network analysis (Hamilton, Ying, and Leskovec 2017), inherently depend on graph-level classification, where label noise can severely compromise the identification of functional groups and the modeling of community behaviors. The lack of attention to graph-level adaptation under noisy labels significantly limits the practical applicability of existing methods in high-impact domains.
@@
In this paper, we investigate the development of an efficient GDA framework for scenarios involving label noise. However, designing such a framework poses several fundamental challenges: (1) Distribution shift undermines loss-based denoising. Conventional noise-robust loss functions are primarily designed for specific domains and often struggle under distribution shifts. In the presence of noisy labels in the source domain, aligning target features with corrupted source representations can lead to noise-aligned embeddings, degrading generalization due to feature misalignment and increased risk of negative transfer. Recent studies in GDA have highlighted that noisy supervision severely hinders feature alignment across domains, especially when relying on pseudo labels or unreliable source signals (Yuan et al. 2023a). These findings underscore the need for noise-aware mechanisms that explicitly account for both label noise and domain discrepancy. (2) Imperfect pseudo labels compromise domain adaptation. Probability-based pseudo-labeling has shown promise in bridging distribution shift and mitigating supervision noise (Yuan et al. 2023a; Yin et al. 2023a). However, the reliability of selected pseudo labels is often compromised by erroneous source annotations, leading to residual noise being transferred into the target domain. In Graph Neural Networks (GNNs), such corrupted pseudo labels can propagate through message passing, triggering self-reinforcing error cascades. As each GNN layer aggregates information from potentially mislabeled neighbors, the accumulated noise progressively deteriorates local neighborhood structures and distorts global representations over successive adaptation rounds (Wang et al. 2024d). (3) Label noise impairs distribution alignment in GDA. Existing methods typically adopt explicit (Long et al. 2015) or implicit (Long et al. 2018) strategies to align feature distributions across domains. However, significant label noise corrupts supervision signals, causing samples to drift toward incorrect class regions and disrupting the formation of domain-invariant features. This misalignment undermines the effectiveness of domain discriminators and hampers reliable adaptation. These challenges call for a unified framework that combines noise-robust representation learning, trustworthy pseudo-label refinement, and alignment strategies that preserve class-level semantics across domains.
@@
Graph Domain Adaptation (GDA) has emerged as a critical research topic, aiming to leverage labeled source domain graphs to enable robust representation learning on unlabeled or sparsely labeled target graphs (Lin et al. 2023; Luo et al. 2023; Liu et al. 2024a). To achieve this, most existing approaches first employ Graph Neural Networks (GNNs) (Kipf and Welling 2017a; Chen et al. 2023) to generate representations for each graph (Wu, Pan, and Zhu 2022; Zhu et al. 2021; Yin et al. 2022b). They then commonly use adversarial learning to implicitly align feature distributions and reduce domain discrepancies, apply pseudo-labeling to iteratively refine predictions in the target domain, or incorporate structure-aware strategies to explicitly align graph-level semantics and topological structures, thereby improving generalization across diverse graph domains (Yin et al. 2023b; Wang et al. 2024b; Liu et al. 2024b). However, these methods often overlook the impact of noisy labels, which can distort learned representations and lead to misaligned distributions and unreliable predictions in the target domain. Although a few label-denoising GDA methods have been proposed, they primarily focus on node-level tasks (Yuan et al. 2023a). To address these limitations, we propose a novel label-denoising domain adaptation method designed for graph-level classification tasks.
Learning with Noise Labels. Learning with noisy labels has garnered significant attention for its crucial role in developing robust models under imperfect supervision, which has been widely used in machine learning and computer vision (Zhu et al. 2024). Existing methods typically address label noise by employing robust loss functions, identifying and filtering out noisy samples, or refining labels through correction mechanisms (Feng et al. 2021; Xu et al. 2025). However, existing methods still insufficiently investigate the interplay between label noise and domain adaptation (Yin et al. 2024a; Zhu et al. 2024). In particular, applying a model trained on the source domain to the target domain can be regarded as a noisy inference process due to distributional shifts inherent in domain adaptation (Yu et al. 2020; Dan et al. 2024). Furthermore, label noise in the source domain can also degrade model performance (Yuan et al. 2023a; Yu et al. 2024). Critically, conventional methods cannot disentangle whether the observed performance degradation is primarily attributable to domain shift or label noise, thereby limiting their ability to address the underlying causes of adaptation failure effectively. To address this challenge, we propose a novel learning framework designed to mitigate the effects of domain shift and label noise.
@@
Datasets. To assess the effectiveness of the proposed NeGPR, we conduct extensive experiments on multiple benchmark datasets from TUDataset, covering diverse types of domain shifts. For structure-based domain shifts, we utilize MUTAGENICITY (Kazius, McGuire, and Bursi 2005), NCI1 (Wale, Watson, and Karypis 2008), FRANKENSTEIN (Orsini, Frasconi, and De Raedt 2015), and PROTEINS (Dobson and Doig 2003), where each dataset is partitioned into source and target domains based on variations in edge, node and graph flux density to simulate structural distribution shifts (Yin et al. 2023b). For feature-based domain shifts, we evaluate NeGPR on PROTEINS, DD, BZR, BZR_MD, COX2, and COX2_MD (Sutherland, O’brien, and Weaver 2003), where domain discrepancies primarily arise from differences in semantic feature distributions. Detailed dataset statistics are provided in Appendix C.
Baselines. We compare the proposed NeGPR with a comprehensive set of competitive baselines on the datasets above. These baselines include two graph kernel methods: WL (Shervashidze et al. 2011) and PathNN (Michel et al. 2023); four general graph neural networks: GCN (Kipf and Welling 2017b), GIN (Xu et al. 2018), GAT (Veličković et al. 2018), and GMT (Baek, Kang, and Hwang 2021); five label denoising methods: Co-teaching (Han et al. 2018), RTGNN (Qian et al. 2023), Taylor-CE (Feng et al. 2021), OMG (Yin et al. 2023c), and SPORT (Yin et al. 2024a); six graph domain adaptation methods: DEAL (Yin et al. 2022b), CoCo (Yin et al. 2023b), SGDA (Qiao et al. 2023), A2GNN (Liu et al. 2024a), StruRW (Liu et al. 2023), and PA-BOTH (Liu et al. 2024b); and two methods that address both label noise and domain adaptation: ALEX (Yuan et al. 2023a) and ROAD (Feng et al. 2023). More detailed descriptions of the baseline settings are provided in Appendix D.
@@ Conclusion
OMG: OMG (Yin et al. 2023c) mitigates label noise in graph classification by jointly optimizing graph embeddings and label reliability. It employs an online sample reweighting mechanism that dynamically adjusts the training focus based on prediction confidence and noise estimation.
SPORT: SPORT (Yin et al. 2024a) addresses label noise by modeling graph classification from a subgraph perspective, identifying reliable substructures through contrastive learning and sample selection. These substructures are then integrated using a noise-tolerant voting mechanism to enhance representation fidelity.
@@
DEAL: DEAL (Yin et al. 2022b) is an unsupervised domain adaptation framework that aligns source and target domains at both the feature and prediction levels using adversarial learning. It further enhances cross-domain generalization by iteratively refining pseudo-labels through entropy minimization and consistency regularization.
CoCo: CoCo (Yin et al. 2023b) introduces a coupled contrastive framework that simultaneously aligns instance-level and class-level representations across domains to capture both fine-grained and semantic-level consistency. It leverages contrastive objectives to enhance feature discrimination and cross-domain coherence.

### v3

@@ Introduction
Graph Domain Adaptation (GDA) (You et al. 2022; Cai et al. 2024) has emerged as a prominent technique for leveraging labeled graph data from a source domain to enhance learning on an unlabeled target graph domain. Its efficacy has been demonstrated across diverse applications, including temporally-evolved social network analysis (Wang et al. 2021), molecular property prediction (Zhu et al. 2023), and protein-protein interaction modeling (Cho, Berger, and Peng 2016). The core paradigm typically involves learning domain-invariant node/graph representations that bridge the distributional shift between source and target domains, thus enabling effective inference on the target data.
@@
However, the success of standard GDA methods crucially relies on the accurately labeled source data. In practice, source domain labels are often corrupted by noise arising from annotation errors (Dai, Aggarwal, and Wang 2021; Yuan et al. 2023b), subjective judgments (Platanios, Dubey, and Mitchell 2016), or inherent ambiguities in data collection (Chen, Shah, and Kyrillidis 2020). This prevalent issue of label noise can severely misguide the learning of domain-invariant representations (Li, Socher, and Hoi 2020), leading to suboptimal or even detrimental adaptation performance on the target domain (Yin et al. 2025). Existing noise label learning methods typically rely on loss function design to mitigate the impact of noisy labels (Han et al. 2018; Natarajan et al. 2013), which selects clean instances for joint training, and robust loss functions (Wei et al. 2020; Li, Socher, and Hoi 2020), which leverage small-loss selection or instance mixture models. While effective in controlled settings, these approaches fall short in the presence of domain shifts. The coexistence of distribution shift and label noise leads to misaligned feature spaces, causing noise-robust losses to erroneously align clean features with noisy targets, thereby amplifying negative transfer (Yu et al. 2020). While recent efforts have been made to address GDA under noisy labels (Yuan et al. 2023a; Wang and Yang 2022), they primarily target node classification tasks, leaving a critical gap in addressing graph-level scenarios. Many real-world applications, such as molecular property prediction (Stokes et al. 2020) and social network analysis (Hamilton, Ying, and Leskovec 2017), inherently depend on graph-level classification, where label noise can severely compromise the identification of functional groups and the modeling of community behaviors. The lack of attention to graph-level adaptation under noisy labels significantly limits the practical applicability of existing methods in high-impact domains.
@@
In this paper, we investigate the development of an efficient GDA framework for scenarios involving label noise. However, designing such a framework poses several fundamental challenges: (1) Distribution shift undermines loss-based denoising. Conventional noise-robust loss functions are primarily designed for specific domains and often struggle under distribution shifts. In the presence of noisy labels in the source domain, aligning target features with corrupted source representations can lead to noise-aligned embeddings, degrading generalization due to feature misalignment and increased risk of negative transfer. Recent studies in GDA have highlighted that noisy supervision severely hinders feature alignment across domains, especially when relying on pseudo labels or unreliable source signals (Yuan et al. 2023a). These findings underscore the need for noise-aware mechanisms that explicitly account for both label noise and domain discrepancy. (2) Imperfect pseudo labels compromise domain adaptation. Probability-based pseudo-labeling has shown promise in bridging distribution shift and mitigating supervision noise (Yuan et al. 2023a; Yin et al. 2023a). However, the reliability of selected pseudo labels is often compromised by erroneous source annotations, leading to residual noise being transferred into the target domain. In Graph Neural Networks (GNNs), such corrupted pseudo labels can propagate through message passing, triggering self-reinforcing error cascades. As each GNN layer aggregates information from potentially mislabeled neighbors, the accumulated noise progressively deteriorates local neighborhood structures and distorts global representations over successive adaptation rounds (Wang et al. 2024b). (3) Label noise impairs distribution alignment in GDA. Existing methods typically adopt explicit (Long et al. 2015) or implicit (Long et al. 2018) strategies to align feature distributions across domains. However, significant label noise corrupts supervision signals, causing samples to drift toward incorrect class regions and disrupting the formation of domain-invariant features. This misalignment undermines the effectiveness of domain discriminators and hampers reliable adaptation. These challenges call for a unified framework that combines noise-robust representation learning, trustworthy pseudo-label refinement, and alignment strategies that preserve class-level semantics across domains.
@@
Graph Domain Adaptation (GDA) has emerged as a critical research topic, aiming to leverage labeled source domain graphs to enable robust representation learning on unlabeled or sparsely labeled target graphs (Lin et al. 2023; Luo et al. 2023; Liu et al. 2024a). To achieve this, most existing approaches first employ Graph Neural Networks (GNNs) (Kipf and Welling 2017a) to generate representations for each graph (Wu, Pan, and Zhu 2022; Zhu et al. 2021; Yin et al. 2022). They then commonly use adversarial learning to implicitly align feature distributions and reduce domain discrepancies, apply pseudo-labeling to iteratively refine predictions in the target domain, or incorporate structure-aware strategies to explicitly align graph-level semantics and topological structures, thereby improving generalization across diverse graph domains (Yin et al. 2023a; Wang et al. 2024a; Liu et al. 2024b). However, these methods often overlook the impact of noisy labels, which can distort learned representations and lead to misaligned distributions and unreliable predictions in the target domain. Although a few label-denoising GDA methods have been proposed, they primarily focus on node-level tasks (Yuan et al. 2023a). To address these limitations, we propose a novel label-denoising domain adaptation method designed for graph-level classification tasks.
Learning with Noise Labels. Learning with noisy labels has garnered significant attention for its crucial role in developing robust models under imperfect supervision, which has been widely used in machine learning and computer vision (Zhu et al. 2024). Existing methods typically address label noise by employing robust loss functions, identifying and filtering out noisy samples, or refining labels through correction mechanisms (Feng et al. 2021; Xu et al. 2025). However, existing methods still insufficiently investigate the interplay between label noise and domain adaptation (Yin et al. 2024; Zhu et al. 2024). In particular, applying a model trained on the source domain to the target domain can be regarded as a noisy inference process due to distributional shifts inherent in domain adaptation (Yu et al. 2020; Dan et al. 2024). Furthermore, label noise in the source domain can also degrade model performance (Yuan et al. 2023a; Yu et al. 2024). Critically, conventional methods cannot disentangle whether the observed performance degradation is primarily attributable to domain shift or label noise, thereby limiting their ability to address the underlying causes of adaptation failure effectively. To address this challenge, we propose a novel learning framework designed to mitigate the effects of domain shift and label noise.
@@
Datasets. To assess the effectiveness of the proposed NeGPR, we conduct extensive experiments on multiple benchmark datasets from TUDataset, covering diverse types of domain shifts. For structure-based domain shifts, we utilize MUTAGENICITY (Kazius, McGuire, and Bursi 2005), NCI1 (Wale, Watson, and Karypis 2008), FRANKENSTEIN (Orsini, Frasconi, and De Raedt 2015), and PROTEINS (Dobson and Doig 2003), where each dataset is partitioned into source and target domains based on variations in edge, node and graph flux density to simulate structural distribution shifts (Yin et al. 2023a). For feature-based domain shifts, we evaluate NeGPR on PROTEINS, DD, BZR, BZR_MD, COX2, and COX2_MD (Sutherland, O’brien, and Weaver 2003), where domain discrepancies primarily arise from differences in semantic feature distributions. Detailed dataset statistics are provided in Appendix C.
Baselines. We compare the proposed NeGPR with a comprehensive set of competitive baselines on the datasets above. These baselines include two graph kernel methods: WL (Shervashidze et al. 2011) and PathNN (Michel et al. 2023); four general graph neural networks: GCN (Kipf and Welling 2017b), GIN (Xu et al. 2018), GAT (Veličković et al. 2018), and GMT (Baek, Kang, and Hwang 2021); five label denoising methods: Co-teaching (Han et al. 2018), RTGNN (Qian et al. 2023), Taylor-CE (Feng et al. 2021), OMG (Yin et al. 2023b), and SPORT (Yin et al. 2024); six graph domain adaptation methods: DEAL (Yin et al. 2022), CoCo (Yin et al. 2023a), SGDA (Qiao et al. 2023), A2GNN (Liu et al. 2024a), StruRW (Liu et al. 2023), and PA-BOTH (Liu et al. 2024b); and two methods that address both label noise and domain adaptation: ALEX (Yuan et al. 2023a) and ROAD (Feng et al. 2023). More detailed descriptions of the baseline settings are provided in Appendix D.
@@ Conclusion
OMG: OMG (Yin et al. 2023b) mitigates label noise in graph classification by jointly optimizing graph embeddings and label reliability. It employs an online sample reweighting mechanism that dynamically adjusts the training focus based on prediction confidence and noise estimation.
SPORT: SPORT (Yin et al. 2024) addresses label noise by modeling graph classification from a subgraph perspective, identifying reliable substructures through contrastive learning and sample selection. These substructures are then integrated using a noise-tolerant voting mechanism to enhance representation fidelity.
@@
DEAL: DEAL (Yin et al. 2022) is an unsupervised domain adaptation framework that aligns source and target domains at both the feature and prediction levels using adversarial learning. It further enhances cross-domain generalization by iteratively refining pseudo-labels through entropy minimization and consistency regularization.
CoCo: CoCo (Yin et al. 2023a) introduces a coupled contrastive framework that simultaneously aligns instance-level and class-level representations across domains to capture both fine-grained and semantic-level consistency. It leverages contrastive objectives to enhance feature discrimination and cross-domain coherence.